assess attitudes towards issues that affect the residents of a​ village, the village randomly chose 800 families to participate in a survey of life attitudes. The village received 628 completed surveys. What is the sample proportion of completed​ surveys? A. 0.886 B. 1.274 C. 0.785 D. 0.628 awnser the correct answer a b c or D thanks
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 628 / 800 = 0.785

0.785 is the sample proportion of completed​ surveys

option C.